OpenStudy (anonymous):

Give an example of a relation that is NOT a function and explain why it is not a function.

OpenStudy (amistre64):

y^2 = x

OpenStudy (amistre64):

i guess 2nd place is silver?

OpenStudy (anonymous):

a function doesn't pass the verticle line test,( for future reference)

OpenStudy (saifoo.khan):

Right.

OpenStudy (amistre64):

verticle line test is an analytic tool :) not proofable

jimthompson5910 (jim_thompson5910):

yeah but it's a good visual tool nonetheless

OpenStudy (amistre64):

for any given input; a function will return only 1 output

OpenStudy (saifoo.khan):

y = x^2 for each value of x, there is, at most, one value for y y = sqrt(x) <=== NOT a function... one value for x = more than one value for y . . . If x = 4 ... then y = + or - 2 <=== multiple values for y

OpenStudy (amistre64):

a circle is a good example of a "not" a function

OpenStudy (amistre64):

y = sqrt(x) IS a function

jimthompson5910 (jim_thompson5910):

y = sqrt(x) is actually a function yet another most popular item of confusion in math...

OpenStudy (anonymous):

\[y ^{2} = x\] Each x has two y values therefore not a function

jimthompson5910 (jim_thompson5910):

y = +-sqrt(x) is not a function

OpenStudy (amistre64):

also we would have to define which variable is the abscessed and ordinated

OpenStudy (anonymous):

y=+-(anything) is not a function :)

OpenStudy (amistre64):

but convention says x is independant and y is dependant

myininaya (myininaya):

a circle with radius bigger than 0

OpenStudy (amistre64):

i drew a circle with a radius smaller than 0 :)

myininaya (myininaya):

lol

OpenStudy (anonymous):

oh.. I thought you drew a circle with radius too big to fit in the draw tool.

jimthompson5910 (jim_thompson5910):

beautiful work of art...wait so you're drawing a blank on this one?

OpenStudy (anonymous):

Can u draw one with an imaginary radius?

OpenStudy (amistre64):

it was either that or make an eddy into another dimension

myininaya (myininaya):

do you the only circle that is a function is a circle with 0 radius

OpenStudy (anonymous):

like an imaginary fried ?

OpenStudy (anonymous):

Give an example of a relation that is NOT a function My sister...

OpenStudy (anonymous):

fried supposed to be friend (freudian slip ?)

OpenStudy (anonymous):

my mother

OpenStudy (anonymous):

Everything u don't know what it is is a relation....